OpenStudy (anonymous):

A ball has an electric charge of +1.5 × 10È© coulombs. At what distance from the ball's center is the electric field strength equal to 5.8 × 10¢ newtons/coulomb? (k = 9.0 × 10© newton·meterÆ/coulombÆ)

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

help

OpenStudy (irishboy123):

use this which gives field strength (E) \[E = \frac{k Q}{r^2}\] \[k = 9 * 10^9\]